Project L requires an initial outlay at t = 0 of $61,000, its expected cash inflows are $14,000 per year for 10 years, and its WACC is 14%. What is the project's payback? Round your answer to two decimal places.
Answer #1

Payback period is the number of years it takes to re-earn the invested amount.

Invested amount = $61,000

Per year cash inflow = $14,000

Cash inflow for 4 years = $14,000 * 4 = $56,000

Amount required in year 5 to complete payback = $61,000 - $56,000 = $5,000

Fraction of year 5 = $5,000/$14,000 = 0.36 year

Payback = 4 + 0.36 years = 4.36 years